MonTEX Diagnostics: System Compatibility Report

Oliver Corff

October 8, 2001

Abstract Running L^{A}TEX 2ε on this article will give information on the status of system software necessary for exploiting advanced features of MonTEX: RL line orientation, vertical text, extended Cyrillic characters for Kazakh and the text-mode gamma.

1 Introduction

MonTEX relies on numerous functions supplied by the existing L^{A}TEX 2ε installation for offering advanced services. Before installing MonTEX it might be interesting to find out which support is offered by the system and which additional packages should be supplied in order to raise MonTEX to its complete functionality.

This document will tell you which functionality is really available on your system and what you should do in case the desired support is missing. Whenever the text informs you that something from CTAN (the Comprehensive TeX Archive Network) should be installed then you have the choice to obtain the required software from CTAN via the Internet or from a CD-ROM with the contents of the CTAN site. In short, this will usually be the same source from which you obtained MonTEX.

2 RL Line Orientation

If you want to be able to print complete paragraphs of Uighur Mongolian texts, and not limit yourself to individual words and phrases, then you will most certainly require the correct writing direction for Uighur Mongolian which starts at the right end of line rather than the left end, as supplied by Implementation Levels III and IV of MonTEX. Apart from its vertical orientation, Uighur Mongolian is a RL writing much like Arab and Hebrew. The RL mechanism is provided by modern L^{A}TEX 2ε installations.

Your system status: During this L^{A}TEX 2ε run, your system did not reveal that it is capable of typesetting RL writings correctly.

Action required: Please make sure you started elatex instead of

!!!

latex for calling e-L^{A}TEX 2ε. If this message persists then you should install a recent TEX distribution from CTAN.

3 Vertical Text

If you want to mix horizontal and vertical text on one page your system needs PostScript support. This implies that your TEX installation can handle insertions of PostScript commands in documents (this is done automatically without your interaction); it does not imply that you can print the results on your printer if your printing equipment is not PostScript-compatible.

Your system status: Your TEX installation can handle vertical text; if your system does not display or print vertical text properly then you should check whether you do have a PostScript printer or interpreter like GhostScript and GhostView.

4 Uighur Mongolian Text Pages

Your system status: Since your system does not offer RL support (see above) this section is meaningless and should be skipped.

5 Extended Cyrillic Support

Beyond the basic Cyrillic support which is provided by MonTEX for writing Modern Xalx Mongolian, Buryat and last but not least, Russian, Kazakh requires additional Cyrillic characters which are not supplied by MonTEX.

Your system status: Since your installation offers the full range of extended Cyrillic characters, it is possible to write Kazakh on your system. No further measures are required.

6 Gamma Support

The Greek gamma γ is used in transliterations for Uighur Mongolian Script. Without support of Modern Greek fonts, only the slim, italic γ used in mathematics is available. If you want to match its typeface with all Latin and Cyrillic letters used in MonTEX then you should make sure the Modern Greek fonts are installed.

Your system status: On this system, the Modern Greek fonts are installed. There is no further action necessary.